# Low Variance Off-policy Evaluation with State-based Importance Sampling

### v4

@@ Abstract
In off-policy reinforcement learning, a behaviour policy performs exploratory interactions with the environment to obtain state-action-reward samples which are then used to learn a target policy that optimises the expected return. This leads to a problem of off-policy evaluation, where one needs to evaluate the target policy from samples collected by the often unrelated behaviour policy. Importance sampling is a traditional statistical technique that is often applied to off-policy evaluation. While importance sampling estimators are unbiased, their variance increases exponentially with the horizon of the decision process due to computing the importance weight as a product of action probability ratios, yielding estimates with low accuracy for domains involving long-term planning. This paper proposes state-based importance sampling, which drops the action probability ratios of sub-trajectories with “negligible states” – roughly speaking, those for which the chosen actions have no impact on the return estimate – from the computation of the importance weight. Theoretical results show this reduces the ordinary importance sampling variance from $O(\exp(H))$ to $O(\exp(X))$ where $X<H$ is the largest subtrajectory with non-negligible states. To identify negligible states, two search algorithms are proposed, one based on covariance testing and one based on state-action values. We formulate state-based variants of ordinary importance sampling, weighted importance sampling, per-decision importance sampling, incremental importance sampling, doubly robust off-policy evaluation, and stationary density ratio estimation. Experiments in four distinct domains show that state-based methods consistently yield reduced variance and improved accuracy compared to their traditional counterparts.
1 Introduction
@@
In reinforcement learning (RL) [18], an agent explores an environment which provides reward signals denoting the desirability of the behaviour, and through this feedback, it learns a policy to optimise a utility function, such as the expected cumulative reward, in that environment. In many domains, including medical interventions (e.g. surgery, diabetes or sepsis), financial decision-making (e.g. advertising or trading), and safe navigation, the exploration process will be too costly as it requires trying out suboptimal policies. In such domains, ideally one would be able to use expert decision-making as an initial starting point and improve the decision-making policy based on a collected data set rather than online interaction with trial-and-error. Such domains motivate the use of off-policy evaluation [25, 7], a class of methods which estimate the utility of the policy to be evaluated using trajectories sampled from a behaviour policy that can be safely used for exploration.
@@
Importance sampling (IS) is a traditional statistical technique which estimates an integral over one distribution of interest while only having samples from another distribution. The expected cumulative reward, discounted or undiscounted, can be seen as such an integral, and therefore IS is an important technique for off-policy evaluation [15, 22, 2]. Such techniques have been used successfully in the context of diabetes treatment [2] and digital marketing [22]. Unfortunately, due to computing the importance weight as a product of action probability ratios across the entire episode, the variance of IS based off-policy evaluation grows exponentially with the horizon $H$ of the decision process, a problem also known as “the curse of horizon” [8]. This makes IS based off-policy evaluation perform more poorly in domains requiring long-term planning (e.g. safe navigation).
To reduce the variance of the importance sampling estimator, Guo et al. (2017) [5] introduce covariance testing, which drops terms (i.e. action probability ratios) from the importance weights based on the expected value of the product of the terms dropped being equal to 1 and the covariance between the estimate and the dropped product being equal to 0. This leads to the implementation of incremental importance sampling (INCRIS), which drops all but the most recent action probability ratios from the importance weight of the per-decision importance sampling (PDIS) estimator [15].
This paper takes a different approach to reduce the variance of importance sampling estimators, which we call state-based importance sampling (SIS). While INCRIS drops all but the $k$ most recent action probability ratios, there is not always an a priori reason why the number $k$ should be related across trajectories. That is, the $i$’th decision in one trajectory is often unrelated to the $i$’th decision in another trajectory – e.g. due to stochasticity of the starting state, the policy, or the transition dynamics – and therefore the $k$ previous time steps may not be the best criterion for dropping action probability ratios. Instead, by dropping action probability ratios associated with a particular state or set of states, the dropped terms for SIS have a related meaning across trajectories. More specifically, SIS drops the action probability ratios corresponding to states that do not affect the estimate when dropped any time they occur in a trajectory; we call such states “negligible states” (see Section 4). We formulate algorithms for identifying negligible states automatically based on covariance testing as well as state-action values (see Section 5). The state-action value based formulation is especially convenient to integrate with other variance reduction techniques, so we demonstrate state-based approaches to a variety of low-variance importance sampling estimators in addition to traditional IS. Beyond ordinary importance sampling, there are also state-based formulation of other importance sampling estimators, doubly robust estimation, and stationary density ratio estimation (see Section 6). Section 7 details the experiments conducted in a deterministic and stochastic lift domain, which include states for which actions do not affect the next state nor the next reward, as well as inventory management and taxi domains as more generic RL benchmarks. Results demonstrate that state-based approaches consistently yield improved accuracy when compared to their traditional counterparts.
2 Problem statement and preliminaries
@@
This present paper is based on finite-horizon undiscounted MDPs, which can be defined based on the tupe $(\mathcal{S},\mathcal{A},r,\mathcal{T})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is a reward function, and $\mathcal{T}:\mathcal{S}\times\mathcal{A}\to\Delta^{\mathcal{S}}$ is the transition dynamics model defined over the probability simplex over the state space, $\Delta^{\mathcal{S}}$. The goal of the reinforcement learning agent is to select actions that maximise the expected cumulative reward, or “return” for short. The decision-making is represented by the policy of the agent, which takes the form $\pi:\mathcal{S}\to\Delta^{\mathcal{A}}$; that is, the policy $\pi$ is a function outputting a vector of action probabilities for a given state.
@@
$\mathcal{G}$ $=\mathbb{E}_{\pi_{e},\mathcal{T}}\left[\sum_{t=1}^{H}r_{t}\right]$
$=\mathbb{E}_{\pi_{b},\mathcal{T}}\left[\sum_{t=1}^{H}r_{t}\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}\right)\right]\,,$
which samples from $\pi_{b}$ and then corrects for the probability under $\pi_{e}$.
A long-standing problem with importance sampling in this manner is that if the horizon is large, then the variance on the estimate of $\mathscr{G}$ is prohibitively large. To see why, define the estimate $\hat{G}$
$\hat{G}$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{H}r_{t}^{(i)}\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}^{(i)}|s_{t}^{(i)})}{\pi_{b}(a_{t}^{(i)}|s_{t}^{(i)})}\right)$
where $i$ indexes the trajectory of the state $s_{t}^{(i)}$, action $a_{t}^{(i)}$, and reward $r_{t}^{(i)}$ at time $t$. Noting that $\hat{G}$ is the sample mean of the random variable $X=\sum_{t=1}^{H}r_{t}\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}\right)$ and assuming $r\in[0,r_{\text{max}}]$, the following provides a general upper bound for the variance of this estimator:
$\text{Var}(\hat{G})$ $=\frac{1}{n}\text{Var}(X)$
@@
$=\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{H}\right)^{2}\,,$
where $\rho_{\text{max}}=\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\pi_{e}(a|s)/\pi_{b}(a|s)$ and the second line follows from assuming rewards in $[0,r_{\text{max}}]$ and applying Popoviciu’s inequality [13], which states that $\text{Var}(X)\leq(M-m)^{2}/4$ for any random variable $X$ with range $[m,M]$. In episodic environments, one may also replace $H$ with the maximal episode length as the effective horizon. In addition to the general upper bound, it is also possible show an exponential lower bound under certain assumptions. For instance, under the assumptions that the Markov chain $\{s_{t},a_{t}\}$ is Harris recurrent, has the geometric drift property, and that $|\log(\rho_{t}|)|<\infty$ for $\rho_{t}=\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}^{(i)}|s_{t})}{\pi_{b}(a_{t}|s_{t})}$, Liu et al. (2020) [9] show that the variance is lower bounded by
$\text{Var}(\hat{G})\geq\frac{G^{2}}{4}\exp\left(\frac{Hc^{2}}{8c_{1}^{2}||B||_{\infty}}\right)-G^{2}\,,$
where $c=\mathbb{E}_{d^{\pi_{b}}}\left[KL(\pi_{b}(a|s)||\pi_{e}(a|s))\right]$ based on the average state-visitation distribution $d^{\pi_{b}}$; $c_{1}$ is a positive constant; and $B$ is the test function of the geometric drift.
3 Related work
In addition to the earlier-mentioned INCRIS algorithm, SIS is related to three lines of work for low-variance importance sampling, namely stationary density ratio estimation, marginalised importance sampling, and conditional importance sampling. The section below briefly discusses these methods followed by a brief comparison to SIS. The section then concludes by referring to other approaches for off-policy evaluation and briefly mentioning important aspects about the application context of off-policy evaluation.
3.1 Approaches related to state-based importance sampling
Stationary density ratio estimation (SDRE) [8] avoids the cumulative product across time and instead is based on the average visitation distribution of state-action pairs. Denoting the average visitation distribution of a state $s_{t}^{(i)}$ as $d_{t}^{\pi}(s_{t}^{(i)})$ and its estimate as $\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})$, SDRE estimates the return as
@@ 3.1 Approaches related to state-based importance sampling
thereby removing the curse of horizon by transforming an exponential dependence on $H$ into a polynomial dependence on $H$. The approach is thereby suitable for large infinite-horizon problems. Unfortunately, the approach requires a suitable model (e.g. a neural network) for estimating the true stationary density ratio. While the approach designs a mini-max loss function to train the model, the accuracy of the approach in estimating the return depends on the accuracy of the model (at worst, the MAE may be equal to the loss). The approach requires assumptions of ergodicity and the state space being finite, and due to requiring a stationary distribution, the approach may not work on finite-horizon MDPs.
Marginalised importance sampling (MIS) [26] is closely related to SDRE. The MIS estimator is given by
@@
where one can note the difference with SDRE by indexing the state distribution $\hat{d}_{t}^{\pi_{e}}$ at a particular time $t$ as well as estimating the reward $\hat{r}_{t}^{\pi_{e}}(s_{t}^{(i)})$ for a particular time $t$. The estimator of the state distribution is updated according to $\hat{d}_{t}^{\pi_{e}}=\hat{\mathcal{T}}_{t}^{\pi_{e}}(s_{t}|s_{t-1})\hat{d}_{t-1}^{\pi_{e}}$, where $\hat{\mathcal{T}}_{t}^{\pi_{e}}(s_{t}|s_{t-1})$ is the empirical importance-weighted visitation of $s_{t-1}\to s_{t}$. The estimator $\hat{r}_{t}^{\pi_{e}}$ is the importance-weighted reward for transitions $s_{t-1}\to s_{t}$. Assumptions are 1) bounded expected value and variance of reward and variance; 2) the behaviour policy has support for all $t$ and all $s_{t}$; and 3) the density ratios and importance weights are bounded. The approach yields a variance that is polynomial in $H$, according to $O(H^{3})$ in general. Looking at specific cases, one can reduce to $O(H^{2})$ when $\pi_{e}\approx\pi_{b}$, and $O(H)$ when the maximal value does not depend on the horizon. Within the formalism of a tabular MDP, i.e. where state space and actions space are limited such that each state-action pair can be frequently visited, Tabular MIS [28] provides an alternative approximation to $\hat{r}_{t}^{\pi_{e}}(s_{t}^{(i)})$ and $\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})$ based on the observed visitation frequencies. Tabular MIS provides a lower bound in line with the Cramer-Rao bound of $\Omega(H^{2}/n)$.
@@
$\hat{G}_{CIS}=\sum_{i=1}^{n}\sum_{t=1}^{H}r_{t}^{(i)}\mathbb{E}[\rho_{1:H}|\phi_{t}]]\,,$ (3)
which yields $\text{Var}(G\mathbb{E}[\rho_{1:H}|\phi_{t}])\leq\text{Var}(G\rho_{1:H})$. Liu et al. (2020) [9] make use of this property to better understand IS, PDIS, and marginalised IS, noting that in this framework the PDIS estimate has $\phi_{t}=s_{1:t},a_{1:t}$ and the marginalised IS estimate has $\phi_{t}=s_{t},a_{t}$. Their results show several findings:
  • For finite horizon, if $\rho_{1:t}$ and $\rho_{1:t}r_{1:t}$ are positively correlated, then PDIS’ variance is lower than traditional importance sampling. The positive correlation happens when the target policy is near-optimal (because then high-density trajectories will be associated with large rewards).
  • For finite horizon, for $t\leq k$, $\text{Cov}(\rho_{1:t}r_{t},\rho_{1:t}r_{k},)\leq\text{Cov}\left(\frac{\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})}{\hat{d}_{t}^{\pi_{b}}(s_{t}^{(i)})}r_{t},\frac{\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})}{\hat{d}_{t}^{\pi_{b}}(s_{t}^{(i)})}r_{k}\right)$, then $\text{Var}(\hat{G}_{PDIS})\leq\text{Var}(\hat{G}_{MIS})$.
  • For infinite horizon, $H\to\infty$, and under some conditions related to ergodicity and recurrence, PDIS has variance $O(H^{2})$ and MIS with oracle importance weights has $O(H^{2})$ but in practice requires approximation of the importance weights, leading to $O(H^{2}(1+\epsilon))$. However, in the general case, these are exponential rather than polynomial.
@@
Rowland et al. (2019) [17] present a related framework for conditional importance sampling with emphasis on trajectories given particular start state-action pair. They introduce novel algorithms arising within the framework, namely return-conditioned, reward-conditioned, and state-conditioned importance sampling. When the equations cannot be solved analytically, the technique often comes with a requirement to form an importance weight regression model, which may increase variance and reduce accuracy.
Compared to the above, SIS estimators can be integrated with a variety of techniques, including ordinary IS, INCRIS, PDIS, and SDRE, which means that it inherits the assumptions of the base-estimator (e.g. with or without function approximation, with stationary distribution assumptions or not). The key property of SIS based estimators is that they work by removing states that are negligible (see Definition 1) from the computation of the importance weight. As such they require a suitable negligible state set as input which is based on either a search algorithm using approximations or prior knowledge. SIS is expected to perform particularly well when one has many states which do not contribute to the expected return and when matched with a base-estimator that is suitable for the domain. The upper bounds for the variance and the MSE for SIS provided in equations 5 and 6 illustrate the effect it can have in case ordinary IS is used as a base-estimator. The integration with other base-estimators is discussed in Section 6, and Section 7 reports experiments comparing state-based variants to their traditional counterparts.
3.2 Other approaches to variance reduction in off-policy evaluation
Other techniques have proven to be useful for off-policy evaluation, and their integration with state-based importance sampling is demonstrated in Section 7. Weighted importance sampling (WIS) and per-decision importance sampling (PDIS) [15] reduce the variance through weighting by normalising based on the sum of trajectory weights and by only accounting for time steps before the decision at the time the action was taken, respectively. The direct method models the reward function which is conceptually simple but requires a model, the accuracy of which often depends on the evaluation policy [3], and any statistical bias in the reward model will be reflected in the estimator. Doubly robust methods [3, 6, 4, 21] use estimators that combine the direct method and importance sampling based on traditional statistical techniques for missing data (e.g. [1]). The doubly robust estimator benefits from the low variance of the direct method and the low bias – or unbiasedness, if the behaviour policy is known – of importance sampling methods. Guo et al. [5] combine INCRIS with options [19], temporally extended actions that help to effectively reduce the horizon of the importance sampling estimator. Rather than competing with these methodologies, state-based methods for importance sampling can be seen as a supplement to these; using the Q-value based identification method (see Section 5), it is straightforward to implement this integration and Section 7 demonstrates this typically reduces the MSE of these approaches.
Other variance techniques not investigated in this paper are also of further interest. Similar to INCRIS, we expect that SIS can also be combined with options for improved accuracy in long-horizon problems. Multiple importance sampling [12] is a traditional Monte Carlo technique that has also been investigated in the context of off-policy evaluation [11]. Estimator selection [25] shares some similarities with state-based importance sampling in the fact that it also selects an estimator by comparing the MSE estimates for different estimators.
3.3 Applying off-policy evaluation

### v5

# 1st David M. Bossens IHPC Agency for Science, Technology and Research Singapore

@@ Abstract
In many domains, the exploration process of reinforcement learning will be too costly as it requires trying out suboptimal policies, resulting in a need for off-policy evaluation, in which a target policy is evaluated based on data collected from a known behaviour policy. In this context, importance sampling estimators provide estimates for the expected return by weighting the trajectory based on the probability ratio of the target policy and the behaviour policy. Unfortunately, such estimators have a high variance and therefore a large mean squared error. This paper proposes state-based importance sampling estimators which reduce the variance by dropping certain states from the computation of the importance weight. To illustrate their applicability, we demonstrate state-based variants of ordinary importance sampling, weighted importance sampling, per-decision importance sampling, incremental importance sampling, doubly robust off-policy evaluation, and stationary density ratio estimation. Experiments in four domains show that state-based methods consistently yield reduced variance and improved accuracy compared to their traditional counterparts.
Index Terms:
off-policy reinforcement learning, off-policy evaluation, importance sampling, variance reduction
I Introduction
@@
In reinforcement learning [1], an agent explores an environment to learn a policy that optimises a utility function, such as the expected cumulative reward, in that environment. In many domains, including medical interventions (e.g. surgery, diabetes or sepsis), financial decision-making (e.g. advertising or trading), and safe navigation, it would be preferable to avoid costly trial-and-error and instead use a known expert or safe policy to gather the data needed for reinforcement learning. Such domains motivate the use of off-policy evaluation [2, 3], a class of methods which estimate the utility of a target policy based on trajectories from a behaviour policy.
@@
Importance sampling (IS) is a traditional statistical technique which estimates an integral over one distribution of interest while only having samples from another distribution. The expected cumulative reward can be seen as such an integral, and therefore IS is an important technique for off-policy evaluation [4, 5, 6]. Such techniques have been used successfully in the context of diabetes treatment [6] and digital marketing [5]. Unfortunately, the variance of IS based off-policy evaluation grows exponentially with the horizon $H$ of the decision process, a problem also known as “the curse of horizon” [7]. This makes IS based off-policy evaluation perform more poorly in domains requiring long-term planning (e.g. safe navigation).
To reduce the variance of off-policy evaluation, this paper proposes a technique that can be used on a wide variety of estimators. The paper makes the following contributions:
  • state-based importance sampling, a class of state-based importance sampling estimators which remove “negligible states” from the importance weight computation;
  • two methods to identify negligible states, one based on covariance testing [8] and one based on state-action values;
  • implementations of state-based variants of ordinary importance sampling, weighted importance sampling [4, 9], per-decision importance sampling [4], incremental importance sampling [8], doubly robust off-policy evaluation [10], and stationary density ratio estimation [7]; and
  • a set of empirical experiments demonstrating the performance of these state-based estimators compared to their traditional counterparts.
II Problem statement and preliminaries
@@
This present paper is based on finite-horizon undiscounted Markov decision processes (MDPs), which can be defined based on the tuple $(\mathcal{S},\mathcal{A},r,\mathcal{T})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is a reward function, and $\mathcal{T}:\mathcal{S}\times\mathcal{A}\to\Delta^{\mathcal{S}}$ is the transition dynamics model defined over the probability simplex over the state space, $\Delta^{\mathcal{S}}$. The goal of the reinforcement learning agent is to select actions that maximise the expected cumulative reward, which is also called the expected return. The decision-making is represented by the policy of the agent, which takes the form $\pi:\mathcal{S}\to\Delta^{\mathcal{A}}$; that is, the policy $\pi$ outputs the action distribution for each given state.
@@
$\mathcal{G}$ $=\mathbb{E}_{\pi_{e}}\left[\sum_{t=1}^{H}r_{t}\right]$
$=\mathbb{E}_{\pi_{b}}\left[\sum_{t=1}^{H}r_{t}\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}\right)\right]\,,$
which samples from $\pi_{b}$ and then corrects for the probability under $\pi_{e}$, using the product $\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}^{(i)}|s_{t}^{(i)})}{\pi_{b}(a_{t}^{(i)}|s_{t}^{(i)})}$ as the importance weight of the trajectory.
A long-standing problem with importance sampling in this manner is that if the horizon is large, then the variance on the estimate of $\mathcal{G}$ is prohibitively large. To see why, define the estimate $\hat{G}_{IS}$
$\hat{G}_{IS}$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{H}r_{t}^{(i)}\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}^{(i)}|s_{t}^{(i)})}{\pi_{b}(a_{t}^{(i)}|s_{t}^{(i)})}\right)\,,$
where $i$ indexes the trajectory of the state $s_{t}^{(i)}$, action $a_{t}^{(i)}$, and reward $r_{t}^{(i)}$ at time $t$. By sample mean of a random variable $X=\sum_{t=1}^{H}r_{t}\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}\right)$ and Popoviciu’s inequality [11],
$\text{Var}(\hat{G}_{IS})$ $=\frac{1}{n}\text{Var}(X)$
@@
$=\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{H}\right)^{2}$
$=\mathcal{O}(\exp(H))\,,$ (1)
where $\rho_{\text{max}}=\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\pi_{e}(a|s)/\pi_{b}(a|s)$ and $[0,r_{\text{max}}]$ is the reward range. In addition to the general upper bound, it is also possible, with some additional assumptions, to show that the variance is lower bounded by $\Omega(\exp(H))$ [12]. Due to the bias-variance decomposition of the mean squared error,
$\text{MSE}(\hat{G}_{IS})=\text{Bias}(\hat{G}_{IS})^{2}+\text{Var}(\hat{G}_{IS})\,,$
where $\text{Bias}(\hat{G}_{IS})=\mathbb{E}_{\pi_{b}}[\hat{G}_{IS}]-\mathcal{G}$, the above results imply an exponential dependency of the MSE on the horizon of the decision problem.
III Related work
Due to the curse of horizon, variance reduction is one of the important problems in off-policy evaluation. We review variance reduction techniques and provide exemplary off-policy evaluation applications.
III-A Variance reduction in importance sampling
While ordinary importance sampling takes a simple average across trajectories, dividing by a denominator $n$, the weighted importance sampling (WIS) technique [4, 9] divides by the sum of the importance weights of the different trajectories. This mitigates the impact of excessively large importance ratios. Its resulting variance converges to 0 for $n\to\infty$ [9] but remains exponential in the horizon [7]. While state-based importance sampling does not work similarly to weighted importance sampling, it shares the commonality that it serves as an easy plug-in to a wide range of estimators.
Another common technique is to remove some of the time steps from the cumulative product, which results in a lower exponent in Eq. II and mitigates the curse of horizon. In per-decision importance sampling (PDIS) [4], one forms different importance weights for each reward $r_{t^{\prime}}$ by relating it to the past time steps until that point, i.e. $\rho_{t^{\prime}}=\prod_{t=1}^{t^{\prime}}\frac{\pi_{e}(a_{t}^{(i)}|s_{t}^{(i)})}{\pi_{b}(a_{t}^{(i)}|s_{t}^{(i)})}$. Incremental importance sampling (INCRIS) [8] drops all but the $k$ most recent action probability ratios from the importance weight of the per-decision importance sampling technique, where the number $k$ is selected based on a covariance test. Removing time steps in this manner is not always a natural solution since the $i$’th decision in one trajectory may not be related to the $i$’th decision in another trajectory. In state-based importance sampling, one drops action probability ratios associated with a particular set of states rather than a particular set of time steps. In doing so, a similar covariance test is proposed as one technique to identify the set of states to drop.
It is also possible to avoid products of action probability ratios in favour of sums over state density ratios. Stationary density ratio estimation (SDRE) [7] avoids the cumulative product across time and instead is based on the average visitation distribution of state-action pairs. Denoting the average visitation distribution of a state $s_{t}^{(i)}$ as $d_{t}^{\pi}(s_{t}^{(i)})$ and its estimate as $\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})$, SDRE estimates the return as
@@ III-A Variance reduction in importance sampling
thereby removing the curse of horizon such that the variance depends polynomially, rather than exponentially, on $H$. The approach is thereby suitable for large and even infinite-horizon problems but requires additional assumptions of ergodicity, a finite state space, and a clearly defined stationary distribution. In similar spirit, Marginalised importance sampling (MIS) [13] is formulated as
@@
where one can note the difference with SDRE by indexing the state distribution $\hat{d}_{t}^{\pi_{e}}$ at a particular time $t$ as well as estimating the reward $\hat{r}_{t}^{\pi_{e}}(s_{t}^{(i)})$ for a particular time $t$. The estimator of the state distribution is updated according to $\hat{d}_{t}^{\pi_{e}}=\hat{\mathcal{T}}_{t}^{\pi_{e}}(s_{t}|s_{t-1})\hat{d}_{t-1}^{\pi_{e}}$, where $\hat{\mathcal{T}}_{t}^{\pi_{e}}(s_{t}|s_{t-1})$ is the empirical importance-weighted visitation of $s_{t-1}\to s_{t}$. The estimator $\hat{r}_{t}^{\pi_{e}}$ is the importance-weighted reward for transitions $s_{t-1}\to s_{t}$. Within the formalism of a tabular MDP, i.e. where state space and actions space are limited such that each state-action pair can be frequently visited, Tabular MIS [14] provides an alternative approximation to $\hat{r}_{t}^{\pi_{e}}(s_{t}^{(i)})$ and $\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})$ based on the observed visitation frequencies. Tabular MIS provides a lower MSE bound in line with the Cramer-Rao bound of $\Omega(H^{2}/n)$. Compared to techniques using state density ratios, which require stationary state distributions and various function approximators, state-based estimators are applicable under varying asssumptions and algorithmic frameworks, depending on which estimator it is plugged into (e.g. ordinary importance sampling vs stationary density ratio estimation).
@@
$\hat{G}_{CIS}=\sum_{i=1}^{n}\sum_{t=1}^{H}r_{t}^{(i)}\mathbb{E}_{\pi_{b}}[\rho_{1:H}|\phi_{t}]]\,,$ (4)
@@
which yields $\text{Var}(G\mathbb{E}_{\pi_{b}}[\rho_{1:H}|\phi_{t}])\leq\text{Var}(G\rho_{1:H})$. Liu et al. (2020) [12] make use of this property to better understand IS, PDIS, and marginalised IS, noting that in this framework the PDIS estimate has $\phi_{t}=s_{1:t},a_{1:t}$ and the marginalised IS estimate has $\phi_{t}=s_{t},a_{t}$. Rowland et al. (2020) [15] present a related framework for conditional importance sampling with emphasis on trajectories given particular start state-action pair. They introduce novel algorithms arising within the framework, namely return-conditioned, reward-conditioned, and state-conditioned importance sampling. When the equations cannot be solved analytically, the technique often comes with a requirement to form an importance weight regression model, which may increase variance and reduce accuracy. Rather than conditioning on states, state-based methods drop selected states from the importance weight every time they occur in the trajectory.
State-based techniques can also be applied when an accurate model is available. In particular, doubly robust (DR) methods [16, 17, 18, 10] use estimators that combine importance sampling with the direct method, which uses an estimated model to infer the return of the evaluation policy, using traditional statistical techniques for missing data (e.g. [19]). The doubly robust estimator benefits from the low variance of the direct method and the low bias – or unbiasedness, if the behaviour policy is known – of importance sampling methods. It is a consistent estimator [10], implying its variance converges to 0 for $n\to\infty$.
Other variance reduction techniques not investigated in this paper are of further interest, including options [20, 8], multiple importance sampling [21, 22], and estimator selection [2].
III-B Applying off-policy evaluation